# The Hidden Link Between RLHF and Contrastive Learning

### v1

@@ Abstract
Alignment of large language models (LLMs) with human values has recently garnered significant attention, with prominent examples including the canonical yet costly Reinforcement Learning from Human Feedback (RLHF) and the simple Direct Preference Optimization (DPO). In this work, we demonstrate that both RLHF and DPO can be interpreted from the perspective of mutual information (MI) maximization, uncovering a profound connection to contrastive learning. Within this framework, both RLHF and DPO can be interpreted as methods that performing contrastive learning based on the positive and negative samples derived from base model, leveraging the Donsker–Varadhan (DV) lower bound on MI (equivalently, the MINE estimator). Such paradigm further illuminates why RLHF may not intrinsically incentivize reasoning capacities in LLMs beyond what is already present in the base model. Building on the perspective, we replace the DV/MINE bound with the Jensen–Shannon (JS) MI estimator and propose the Mutual Information Optimization (MIO). Comprehensive theoretical analysis and extensive empirical evaluations demonstrate that MIO mitigates the late-stage decline in chosen-likelihood observed in DPO, achieving competitive or superior performance across various challenging reasoning and mathematical benchmarks. The code is available at: https://anonymous.4open.science/r/MIO-63E6/
@@ 1 Introduction
Recently, large language models (LLMs) have attained state-of-the-art performance on a wide array of natural-language tasks [1]. Aligning their outputs with human value is commonly achieved through the methodology of reinforcement learning from human feedback (RLHF) [2][3], whose fine-tuning recipe has become one of the industry standards. However, its dependence on two step reinforcement learning presents challenges [4], such as computational inefficiency and training instability. To mitigate these limitations, alternative one-step approaches such as direct preference optimization (DPO) [5] and its variants [6][7] have been proposed , which has successfully eliminated the need for explicit reward modeling. Specifically, an implicit reward based on the likelihood of preference data is defined, which results in significant gains in efficiency while preserving competitive performance.
Major Questions 1. Could RLHF be theoretically framed as a form of contrastive learning? 2. Does RLHF fundamentally extend a base model’s capabilities or only amplify existing ones? 3. How can we overcome DPO’s limitations to further improve the reasoning and mathematical performance of fine-tuned LLMs?
@@
Several recent works have hinted at conceptual links between RLHF and imitation learning [8]. DIL [4] show that optimizing a policy with RLHF implicitly maximizes the likelihood of human-preferred outputs, formally proves that RLHF for aligning language models could be view as an imitation learning problem. IRL[9] examines the pretraining and fine-tuning of LLMs from an imitation learning standpoint, argue that standard next-token fine-tuining is a form of imitation learning. Beyond imitation learning, several works have drawn intuitive parallels between Direct Preference Optimization (DPO) and contrastive learning [10]. For example, Xu et al. [11] and Chen et al. [12] treat DPO as a form of contrastive learning; however, their analysis still lacks a formal theoretical foundation. Such observation motivates our central research question: Is there a unified theoretical framework that rigorously establishes the connection between RLHF and contrastive learning?
Figure 1: RLHF performs as contrastive learning: PPO/DPO pushes the base model toward the positive model $\pi_{\text{chosen}}$ and away from the negative model $\pi_{\text{rejection}}$. Search trees are generated by repeatedly sampling from the models for a given problem. The gray indicates paths that are unlikely to be sampled by the model, while the black indicates paths that are likely to be sampled. “+1” indicates the correct paths, which have positive rewards. The base model contains both the correct and false path. The positive model is more likely to reach the correct path, whereas the negative model is more likely to follow incorrect paths. After the RLHF process, the policy model becomes closer to the positive model and diverges from the negative model.
Recently, it is widely believed that Reinforcement Learning with Variable Rewards (RLVR), a variant of RLHF, facilitates the continual self-improvement of LLMs, thereby enabling them to acquire novel reasoning capabilities beyond those of their corresponding base models [13]. However, recent empirical findings challenge this assumption. Specifically, Yue et al. [14] reports that RLVR training does not, in fact, induce fundamentally new reasoning trajectories. Their analysis reveals that the reasoning paths exhibited by RL-trained models largely fall within the sampling distribution of the base models, implying that the reasoning capabilities demonstrated post-RL training may already be latent in the base models prior to fine-tuning. Nevertheless, the observed phenomenon in this work lacks a rigorous theoretical explanation. This brings us to our second research question: Is there a theoretical framework that can formally establish whether RLHF truly extends the capabilities of a base model, or merely amplifies its pre-existing capacities?
@@
RLHF and its variants fundamentally adhere to a reward maximization objective, often instantiated through parametric models such as the Bradley–Terry (BT) model [15]. This modeling paradigm, while effective in aligning with preference data, has been shown to suffer from overfitting [16, 17], frequently leading to suboptimal generalization and misalignment with true user preferences [18, 19]. Recent empirical studies further indicate that DPO and its variants progressively focus on unlearning the rejected responses. This process inadvertently increases the model’s tendency to generate out-of-distribution responses, rather than reinforcing alignment with the chosen responses [18]. From a theoretical perspective, Yan et al. [20] analyzes the optimization dynamics of DPO and demonstrates that the shared tokens between chosen and rejected responses lead to gradient interference. This interference introduces instability and a coupled degradation, whereby suppressing rejected responses inadvertently reduces the model’s confidence in the selected responses. Such simultaneous unlearning of both rejected and chosen outputs ultimately degrades downstream performance—particularly in reasoning-intensive and mathematical tasks [17, 16, 6]. This observation motivates our third research question: Can we design a principled algorithm that overcomes inherent shortcomings of DPO, while effectively enhancing model performance on complex reasoning and mathematical tasks?
This paper makes the following contributions: (1) We present a unified mutual-information perspective on RLHF and contrastive learning, demonstrating that both RLHF and DPO optimize a Donsker–Varadhan (DV) [21, 22] lower bound on mutual information, thereby aligning with contrastive objectives. (2) This framework provides a theoretical explanation for why RLHF fails to elicit novel reasoning trajectories: it can only amplify reasoning patterns already present in the base model. (3) To overcome the degradation issue in DPO, we propose the Mutual Information Optimisation (MIO), a one-step algorithm based on a Jensen–Shannon MI estimator [23], which achieves state-of-the-art performance on multiple reasoning and mathematical benchmarks. (4) We further provide a rigorous theoretical analysis showing that MIO preserves meaningful learning signals on chosen responses, even as rejected responses are suppressed.
@@ 2 Rethinking RLHF: A Contrastive Learning Perspective
Given a prompt $x=[\text{x}_{1},\text{x}_{2},...]$, we are presented with two candidate responses: $y_{w}=[\text{y}_{1},\text{y}_{2},...]$, the response preferred by humans; and $y_{l}$, the response with lower preference. Our goal is to fine-tuning the language model, enhancing its propensity to generate responses that are consistent with human preferences. Firstly, we assume that all positive examples $y_{w}$ are drawn from a preferred language model $\pi_{\text{chosen}}$, while negative examples $y_{l}$ are drawn from a dispreferred language model $\pi_{\text{rejection}}$. We assume that both $\pi_{\text{chosen}}$ and $\pi_{\text{rejection}}$ are energy-based models (EBMs) [24], as discussed in Appendix A. Our objective is to adjust the target model $\pi_{\theta}$ to maximize mutual information with ${\pi}_{\text{chosen}}$ while minimizing mutual information with $\pi_{\text{rejection}}$.
@@
In Appendix B, we show that the contrastive learning objective and the imitation learning objective are equivalent, since increasing one objective decreases the other.
Since mutual information (MI) cannot be directly computed, we introduce an estimator for the lower bound of mutual information, namely the Mutual Information Neural Estimation (MINE) [22]:
@@
where $\bar{\pi}(\cdot|x)=\frac{1}{2}\pi_{\text{chosen}}(\cdot|x)+\frac{1}{2}\pi_{\text{rejection}}(\cdot|x)$. Considering the following Monte Carlo estimator:
@@
We substitute Equation (6) into Equation (5), can estimate the lower bound of mutual information as:
$\max_{\theta}\hat{I}=\log\left(\frac{e^{T_{\phi}(\pi_{\theta},\pi_{\text{chosen}})}}{\frac{1}{M}\sum_{i=1}^{M}e^{T_{\phi}(\pi_{\theta},\pi_{\text{chosen}})}+\frac{1}{N}\sum_{i=1}^{N}e^{T_{\phi}(\pi_{\theta},\pi_{\text{rejection}})}}\right).$ (7)
@@
$\max_{\theta}\log\sigma\left(T_{\phi}(\pi_{\theta},\pi_{\text{chosen}})-T_{\phi}(\pi_{\theta},\pi_{\text{rejection}})\right),$ (8)
@@
Furthermore, if we select $T_{\phi}(\pi_{\theta},\pi_{\text{chosen}})=\log\frac{\pi_{\theta}}{\pi_{\text{chosen}}}$, and substitute it into Equation (8), we can arrive at the objective function of Direct Preference Optimization (DPO) [5]:
@@
Keep $T_{\phi}(\pi_{\theta},\pi)=\log\frac{\pi_{\theta}}{\pi}$. Only substituting this into the right-hand side of the MINE objective (4) and applying certain approximations yields the following objective used in the second stage of RLHF, where the policy model is updated using the reward model[19, 25]:
@@
The approximation error introduced here is upper-bounded (see Appendix D), and the bound depends on the ratio $\frac{\pi_{\theta}(\cdot\mid x)}{\pi_{\text{ref}}(\cdot\mid x)}$: the smaller this ratio, the smaller the approximation error. This also explains why PPO introduces a clipping function to constrain the update magnitude of the policy—large deviations in this ratio lead to large approximation errors, which in practice degrade performance. As the derivatives of $\hat{I}(\pi_{\theta},\pi_{\text{chosen}})$ and $\hat{I}(\pi_{\theta},\pi_{\text{rejection}})$ are opposites, maximizing $\hat{I}(\pi_{\text{chosen}},\pi_{\theta})$ simultaneously minimizes $\hat{I}(\pi_{\text{rejection}},\pi_{\theta})$. This shows that optimizing the mutual information between $\pi_{\theta}$ and $\pi_{\text{chosen}}$ inherently minimizes the mutual information between $\pi_{\theta}$ and $\pi_{\text{rejection}}$, thereby transforming the task into a contrastive learning framework.
@@
Furthermore, through the selection of $T_{\phi}(\pi_{\theta},\pi_{\text{chosen}})=\log\frac{\pi_{\theta}}{\pi_{\text{chosen}}}$, we can draw some insights:
@@
    After learning the reward model in RLHF, the subsequent update of the policy model is equivalent to a contrastive learning task. This phase can be seen as maximizing the mutual information between $\pi_{\theta}$ and the positive sample $\pi_{\text{chosen}}$, while minimizing the mutual information with the negative sample $\pi_{\text{rejection}}$. If we limit the search for $T$ to a subset $\mathcal{F}_{\text{sub}}$, (see Appendix C.4), we can obtain the DPO loss function.
Answer to Question 2: RLHF does not fundamentally extend a base model’s capabilities.
Definition 2.1 (Model Capacity).
Given an input question $x$ and its correct answer $y_{w}$, we say that a model $\pi_{\theta}$ has the capacity to correctly answer the question if $\pi_{\theta}(y_{w}\mid x)\neq 0$; that is, the model assigns non-zero probability to the correct answer.
Corollary 2.1.
RLHF does not inherently incentivize the emergence of reasoning capabilities in large language models that go beyond those already present in the base model.
In practice, the base model is typically frozen and used as a reference model. In this setting, positive samples are drawn from $\pi_{\text{chosen}}$, which can be interpreted as an enhanced version of the base model, while negative samples represent a degraded variant. Consequently, the optimal policy $\pi_{\theta}$ is expected to fully align with the $\pi_{\text{chosen}}$. However, since $\pi_{\text{chosen}}$ is defined as an energy-based reweighting of $\pi_{\text{reference}}$, its support is inherently limited by that of the base model. If the base model assigns zero probability to a correct reasoning path (i.e., $\pi_{\text{reference}}(y_{w}\mid x)=0$), then even a well-trained reward model cannot recover or induce such reasoning capability in $\pi_{\theta}$.
RLHF does not fundamentally extend a base model’s capabilities; rather, it amplifies the reasoning abilities already present within base model. It refines and reinforces these existing patterns but does not create entirely new capabilities. This provides a theoretical explanation for experimental findings in some studies [14], which observe that RLHF does not enhance the capabilities of base model.
3 Beyond DPO: From DV to JSD Mutual-Information Optimisation
3.1 Why should JSD be used instead of MINE in Alignment?
Subsequent research has shown that the DV-based objective used in MINE suffers from significant instability in practice [26]. Specifically, its variance can grow exponentially with the true mutual information (MI) [27], leading to exploding gradients and requiring very large batch sizes to maintain stable training [28]. Poole et al. [26] further demonstrate that DV-style bounds like MINE exhibit low bias but extremely high variance, causing erratic gradient updates.
In the work [29], it is emphasized that accurate MI estimation is intrinsically challenging: high bias or high variance can lead to overfitting of spurious correlations. Thus, in contrastive learning, stable MI lower bounds (e.g., MINE or the Jensen–Shannon (JS) bound) should be prioritized over maximizing numerical MI estimates.
Additionally, Sordoni et al. [30] point out that the MINE bound is upper-bounded by $\log K$, where $K$ is the number of negative samples. When the true MI greatly exceeds $\log K$, the bound severely underestimates MI.
Collectively, these studies [28, 29] highlight that in regimes with few negative samples, MINE suffers from prohibitively large variance [26], leading to unstable gradients. Therefore, replacing MINE with a variance-reduced alternative, such as the JSD estimator [31], is crucial for reliable optimization.

### v2

@@ Abstract
Alignment of large language models (LLMs) with human values has recently garnered significant attention, with prominent examples including the canonical yet costly Reinforcement Learning from Human Feedback (RLHF) and the simple Direct Preference Optimization (DPO). In this work, we demonstrate that both RLHF and DPO can be interpreted from the perspective of mutual information (MI) maximization, uncovering a profound connection to contrastive learning. Within this framework, both RLHF and DPO can be interpreted as methods that performing contrastive learning based on the positive and negative samples derived from base model, leveraging the Donsker–Varadhan (DV) lower bound on MI (equivalently, the MINE estimator). Such paradigm further reveals why RLHF cannot correctly update when the probability of sampling the correct answer from the base model is close to 0, which often leads to RLHF failing to incentivize reasoning capacities in LLMs beyond what is already present in the base model. Building on the perspective, we replace the DV/MINE bound with the Jensen–Shannon (JS) MI estimator and propose the Mutual Information Optimization (MIO). Comprehensive theoretical analysis and extensive empirical evaluations demonstrate that MIO mitigates the late‑stage decline in chosen-likelihood observed in DPO, achieving competitive or superior performance across various challenging reasoning and mathematical benchmarks. The code is available at: https://anonymous.4open.science/r/MIO-63E6/
@@ 1 Introduction
Recently, large language models (LLMs) have attained state-of-the-art performance on a wide array of natural-language tasks (Bai et al., 2022). Aligning their outputs with human value is commonly achieved through the methodology of reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022)(Stiennon et al., 2022), whose fine-tuning recipe has become one of the industry standards. However, its dependence on two step reinforcement learning presents challenges (Xiao et al., 2025), such as computational inefficiency and training instability. To mitigate these limitations, alternative one-step approaches such as direct preference optimization (DPO) (Rafailov et al., 2023) and its variants (Meng et al., 2024)(Tajwar et al., 2024) have been proposed , which has successfully eliminated the need for explicit reward modeling. Specifically, an implicit reward based on the likelihood of preference data is defined, which results in significant gains in efficiency while preserving competitive performance.
Major Questions 1. Could RLHF be theoretically framed as a form of contrastive learning?
2. Why do PPO/DPO fail to update when $\pi_{\theta}(y^{*}|x^{*})$ is too small?
3. How can we overcome DPO’s limitations to further improve the reasoning and mathematical performance of fine-tuned LLMs?
@@
Several recent works have hinted at conceptual links between RLHF and imitation learning (Ho & Ermon, 2016). DIL (Xiao et al., 2025) show that optimizing a policy with RLHF implicitly maximizes the likelihood of human-preferred outputs, formally proves that RLHF for aligning language models could be view as an imitation learning problem. IRL(Wulfmeier et al., 2024) examines the pretraining and fine-tuning of LLMs from an imitation learning standpoint, argue that standard next-token fine-tuining is a form of imitation learning. Beyond imitation learning, several works have drawn intuitive parallels between Direct Preference Optimization (DPO) and contrastive learning (Jaiswal et al., 2021). For example, Xu et al. (2024a) and Chen et al. (2024) treat DPO as a form of contrastive learning; However, these works lack theoretical depth and their analysis still lacks a formal theoretical foundation. Such observation motivates our central research question: Is there a unified theoretical framework that rigorously establishes the connection between RLHF and contrastive learning?
Recent empirical studies have challenged the widely held belief that reinforcement learning with variable rewards (RLVR) or standard KL-regularized RLHF can continuously unlock genuinely novel reasoning trajectories. Yue et al. (2025) observe that, on benchmarks such as GSM8K and MATH, reasoning chains produced after RLHF almost always fall within the sampling support of the base model. This suggests that RL primarily amplifies pre-existing capabilities rather than discovering new ones. However, the underlying causes of this phenomenon remain unclear. Additionally, some studies have pointed out that when the initial value of $\pi_{\theta}(y_{l}\mid x)$ is too small, DPO training tends to collapse. This implies that, whether for positive responses $y_{w}$ or negative responses $y_{l}$, if their probabilities become too small, RLHF ceases to be effective. This leads to our second question: Why does RLHF (PPO/DPO) lose its ability to update when the probability of a specific answer $y^{*}$ sampled by the base model is too small? This paper provides a rigorous theoretical explanation for these issues, which can ultimately be attributed to the Donsker–Varadhan mutual information estimator $I_{DV}$.
@@
RLHF and its variants fundamentally adhere to a reward maximization objective, often instantiated through parametric models such as the Bradley–Terry (BT) model (Bradley & Terry, 1952). This modeling paradigm, while effective in aligning with preference data, has been shown to suffer from overfitting (Yuan et al., 2024; Pal et al., 2024), frequently leading to suboptimal generalization and misalignment with true user preferences (Xu et al., 2024b; Wu et al., 2024). Recent empirical studies further indicate that DPO and its variants progressively focus on unlearning the rejected responses. This process inadvertently increases the model’s tendency to generate out-of-distribution responses, rather than reinforcing alignment with the chosen responses (Xu et al., 2024b). From a theoretical perspective, Yan et al. (2025) analyzes the optimization dynamics of DPO and demonstrates that the shared tokens between chosen and rejected responses lead to gradient interference. This interference introduces instability and a coupled degradation, whereby suppressing rejected responses inadvertently reduces the model’s confidence in the selected responses. Such simultaneous unlearning of both rejected and chosen outputs ultimately degrades downstream performance—particularly in reasoning-intensive and mathematical tasks (Pal et al., 2024; Yuan et al., 2024; Meng et al., 2024). This observation motivates our third research question: Can we design a principled algorithm that overcomes inherent shortcomings of DPO, while effectively enhancing model performance on complex reasoning and mathematical tasks?
This paper makes the following contributions: (1) We present a unified mutual-information perspective on RLHF and contrastive learning, demonstrating that both RLHF and DPO optimize a Donsker–Varadhan (DV) (Donsker & Varadhan, 1983; Belghazi et al., 2018a) lower bound on mutual information, thereby aligning with contrastive objectives. (2) We provides a theoretical explanation for why RLHF cannot stimulate new reasoning paths in models, and why DPO fails to update effectively when $\pi_{\theta}(y\mid x)$ is too small: this is attributed to $I_{DV}$, which fails to provide the correct training signal to the policy model when $\pi_{\theta}(y\mid x)$ is too small. (3) Replaces the DV mutual information estimator with the JS mutual information estimator(Hjelm et al., 2019), resulting in a new method, MIO, which avoids the issue of synchronous descent and performs well on eight reasoning and mathematics-related problems. The paper rigorously proves that MIO does not encounter the problem of synchronous descent.
@@ 2.1 RLHF is a form of Contrastive Learning
Given a prompt $x=[\text{x}_{1},\text{x}_{2},$...], we are presented with two candidate responses: $y_{w}=[\text{y}_{1},\text{y}_{2},$...], the response preferred by humans; and $y_{l}$, the response with lower preference. Our goal is to fine-tuning the language model, enhancing its propensity to generate responses that are consistent with human preferences. Firstly, we assume that all positive examples $y_{w}$ are drawn from a preferred language model $\pi_{\text{chosen}}$, while negative examples $y_{l}$ are drawn from a dispreferred language model $\pi_{\text{rejection}}$. We assume that both $\pi_{\text{chosen}}$ and $\pi_{\text{rejection}}$ are energy-based models (EBMs)  (Lecun et al., 2006), as discussed in Appendix A. Our objective is to adjust the target model $\pi_{\theta}$ to maximize mutual information with ${\pi}_{\text{chosen}}$ while minimizing mutual information with $\pi_{\text{rejection}}$.
@@
In Appendix B, we demonstrate that the contrastive learning and imitation learning objectives are equivalent, as improving one objective decreases the other and the strict local maximum of the former corresponding to the strict local minimum of the latter.
Since mutual information (MI) cannot be directly computed, we introduce an estimator for the lower bound of mutual information, namely the Mutual Information Neural Estimation (MINE) Belghazi et al. (2018a):
@@
where $\bar{\pi}(\cdot|x)=\frac{1}{2}\pi_{\text{chosen}}(\cdot|x)+\frac{1}{2}\pi_{\text{rejection}}(\cdot|x)$.The right-hand side of equation equation 5 is also referred to as $I_{DV}$  (Hjelm et al., 2019). Considering the following Monte Carlo estimator:
@@
We substitute  equation 6 into  equation 5, can estimate the lower bound of mutual information as:
$\max_{\phi}\hat{I}=\log\left(\frac{e^{T_{\phi}(\pi_{\theta},\pi_{\text{chosen}})}}{\frac{1}{M}\sum_{i=1}^{M}e^{T_{\phi}(\pi_{\theta},\pi_{\text{chosen}})}+\frac{1}{N}\sum_{i=1}^{N}e^{T_{\phi}(\pi_{\theta},\pi_{\text{rejection}})}}\right).$ (7)
@@
$\max_{\phi}\log\sigma\left(T_{\phi}(\pi_{\theta},\pi_{\text{chosen}})-T_{\phi}(\pi_{\theta},\pi_{\text{rejection}})\right),$ (8)
@@
Furthermore, instead of retraining a separate neural network $T_{\phi}$, we constrain the T to the log-ratio sub-family (see in Appendix C.2):
$T(\pi_{\theta},\pi_{chosen})=\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\mathrm{chosen}}(y_{w}|x)},\qquad T(\pi_{\theta},\pi_{rejection})=\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\mathrm{rejection}}(y_{l}|x)}.$ (9)
and substitute it into  equation 8, we can arrive at the objective function of Direct Preference Optimization (DPO)(Rafailov et al., 2023):
@@
Keep constrain the T to the log-ratio sub-family equation 9. Only substituting this into the right-hand side of the MINE objective equation 4 and applying certain approximations yields the following objective used in the second stage of RLHF, where the policy model is updated using the reward model(Wu et al., 2024; Ghosh et al., 2020):
@@
The approximation error introduced here is upper-bounded (see Appendix E), and the bound depends on the ratio $\frac{\pi_{\theta}(\cdot\mid x)}{\pi_{\text{ref}}(\cdot\mid x)}$: the smaller this ratio, the smaller the approximation error. This offers us a new and insightful perspective about why PPO introduces a clipping function to constrain the update magnitude of the policy—large deviations in this ratio lead to large approximation errors, which in practice degrade performance. As the derivatives of $\hat{I}(\pi_{\theta},\pi_{\text{chosen}})$ and $\hat{I}(\pi_{\theta},\pi_{\text{rejection}})$ are opposites, maximizing $\hat{I}(\pi_{\text{chosen}},\pi_{\theta})$ simultaneously minimizes $\hat{I}(\pi_{\text{rejection}},\pi_{\theta})$. This shows that optimizing the mutual information between $\pi_{\theta}$ and $\pi_{\text{chosen}}$ inherently minimizes the mutual information between $\pi_{\theta}$ and $\pi_{\text{rejection}}$, thereby transforming the task into a contrastive learning framework.
@@
Furthermore, we can draw some insights:
@@
[…]
2.2 The DV mutual information estimator causes RLHF collapse.
Answer to Question 2: Why do PPO/DPO fail to update when $\pi_{\theta}(y^{*}|x^{*})$ is too small? A mutual information explanation
In essence, based on our analysis, both RLHF and DPO can be viewed as approximate procedures for optimizing a DV/MINE-based mutual information estimator $I_{\mathrm{DV}}(\theta)$. However, such estimators fail to provide meaningful gradient directions for a specific response $y^{\star}$ when the reference policy $\pi_{\mathrm{ref}}(y^{\star}\mid x^{\star})$ approaches zero. Since $\pi_{\theta}$ is typically initialized as a frozen copy of the reference model, if the initial $\pi_{\theta}(y^{\star}\mid x^{\star})$ is close to zero, then no matter how well the reward model is trained, using it to further optimize $\pi_{\theta}$ becomes ineffective. The root cause of this phenomenon is precisely the DV/MINE estimator $I_{\mathrm{DV}}(\theta)$. To formalize this claim, we present two theorems:
[…]
$\big\langle\nabla_{\theta}I_{\mathrm{DV}}(\theta),\,\nabla_{\theta}\log\pi_{\theta}(y^{\star}\mid x^{\star})\big\rangle=0.$
This explains why PPO-style methods (e.g., GRPO) typically require a cold start stage prior to increase the probability of the base model sampling the correct answer $y^{\star}$: when the probability of sampling the correct answer $y^{\star}$ is too small, the cold start could increase it to prevent policy gradients from vanishing due to $\pi_{\mathrm{ref}}(y^{\star}\mid x^{\star})=0$.
Some readers may object that, in practice, neural networks assign nonzero probability to every sequence (albeit extremely small), so Theorem 1 may not hold exactly. To address this, we provide a strengthened result:
Theorem 2 (DV/MINE Starvation Theorem Pro). As $\pi_{\theta}(y^{\star}\mid x^{\star})$ approaches zero, the DV/MINE estimator $I_{\mathrm{DV}}(\theta)$ yields increasingly misleading gradients. Formally,
$\big|\big\langle\nabla_{\theta}I_{\mathrm{DV}}(\theta),\,\nabla_{\theta}\log\pi_{\theta}(y^{\star}\mid x^{\star})\big\rangle\big|\;\leq\;2L\,\pi_{\theta}(y^{\star}\mid x^{\star}),$
where $L$ is a constant. A full proof is provided in Appendix  H.
This theorem offers a mutual-information perspective on why off-policy settings in PPO/GRPO often lead to degraded training performance: under an off-policy regime, the probability $\pi_{\theta}(y\mid x)$ of sampling a trajectory is extremely small, which makes the gradient of $I_{\mathrm{DV}}$ nearly orthogonal to the policy gradient. As a result, the mutual information estimator cannot be used to approximate the gradient of the policy model $\pi_{\theta}$. This also provides an intuitive insight: the effectiveness of DV/MINE-based gradient estimation is highly sensitive to the initial values of $\pi_{\theta}(y\mid x)$. Once these probabilities are too small, the estimator $I_{\mathrm{DV}}$ becomes too hungry to update.
In summary, the DV/MINE estimator introduces serious drawbacks: not only does it induce large gradient variance (see in Appendix D), but it also leads to policy-gradient starvation whenever $\pi_{\mathrm{ref}}(y\mid x)$ is near zero. This severely hinders the learning dynamics of RLHF. Mitigation strategies such as cold-start initialization are often used to increase $\pi_{\mathrm{ref}}(y\mid x)$, thereby reducing gradient starvation. However, the fundamental culprit remains the DV/MINE estimator itself. This motivates an important research question: are there effective alternatives to DV/MINE-based mutual information estimators?
3 Methods: Mutual-Information Optimisation
3.1 Let us replace DV/MINE Estimator with JSD in alignment.